# Measurement as Bricolage: Examining How Data Scientists Construct Target Variables for Predictive Modeling Tasks

### v1

@@ 2.5. Measurement and Validity in AI
A growing line of research traces AI system failures to validity pitfalls (Jacobs and Wallach, 2021; Milli et al., 2021; Stray et al., 2021; Guerdan et al., 2023b; Goyal et al., 2022; Obermeyer et al., 2019; Mullainathan and Obermeyer, 2021; Coston et al., 2023). In the quantitative social sciences, the validity of a measurement instrument is established via a multifaceted assessment along several dimensions. Construct validity describes the extent to which a measurement instrument wholly and fully reflects the theoretical construct it purports to measure (Jacobs and Wallach, 2021). In predictive modeling, construct validity is most often discussed in connection to whether the target variable predicted by a model reflects a theoretical construct of interest to model developers (Guerdan et al., 2023b; Coston et al., 2023).  [^1] Jacobs and Wallach (2021) provide a detailed review of sub-dimensions of construct validity as they pertain to measuring fairness properties of algorithmic systems. Internal and external validity are also important dimensions in experimental research. Internal validity describes the extent to which an experimental study can confidently establish a causal relationship between independent and dependent variables (Gergle and Tan, 2014). In predictive modeling, it relates to the existence of a defensible causal relationship between features and a target variable (Coston et al., 2023). External validity, on the other hand, measures how well an analysis can be generalized beyond the specific conditions of the study. This dimension has been linked to the extent to which a predictive model’s in-distribution performance generalizes across real-world deployment contexts (Coston et al., 2023).
@@
We expand upon Mussgnug (2022)’s argument by characterizing how data scientists weigh validity and predictive performance during problem formulation. In support of Mussgnug (2022)’s argument, we find that many prediction tasks pursued by data scientists are indeed measurement problems. For example, some participants in our study used supervised learning models to measure concepts recorded at future points in time — e.g., “10-year cardiovascular risk.” Others used supervised learning models as measurement instruments while measuring a concept at a current point of time — e.g., to score the “authenticity” of a student’s writing. In both cases, data scientists heavily anchored on predictive performance while evaluating their problem formulation. However, counter to Mussgnug (2022)’s argument, we find that data scientists do also engage with validity concepts during problem formulation, albeit with a more pragmatic, situated approach that lacks formal terminology. We uncover how data scientists evaluate validity in our findings ($\S$§ 4), and unpack implications for data science pedagogy ($\S$§ 5.2.3) and tooling support ($\S$§ 5.3.4) in the discussion.
@@ 3. Methods
To examine how data scientists construct target variables for predictive modeling tasks, we interviewed 15 data scientists from the education (N=8) and healthcare (N=7) sectors who had prior experience with developing and evaluating predictive models. We conducted one-hour semi-structured interviews to understand participants’ current practices and perceived challenges for measurement as they engaged in problem formulation.
@@ 3.1.1. Directed Storytelling
During the first interview segment, we used a directed storytelling approach (Evenson, 2016), inviting data scientists to reflect on a time in the past when they developed and evaluated a predictive model. We centered questions in this segment around data scientists’ target variable construction process, a component of problem formulation which prior literature has shown is both complex and consequential (Obermeyer et al., 2019).
To ground our discussions, we first asked participants to share a few concrete experiences where they had developed and/or evaluated predictive models as part of their role. Building on these concrete experiences, the interviewer then narrowed in on one of the projects mentioned and asked participants the following follow-up questions, in a semi-structured fashion:
@@
We also asked participants a range of additional follow-up questions when these were not already covered by participants during their earlier responses, such as measurement guidelines and best practices they consulted throughout their project. On average, the first segment lasted 30-40 minutes of the interview.
@@ 3.1.2. Vignette-Based Problem Formulation Task
During the second segment, we asked participants to reason about a hypothetical problem formulation as we walked them through a vignette. We aimed to better understand how participants reason about measurement decisions on-the-fly, while potentially surfacing challenges and opportunities they might not have recalled while reflecting on their prior experiences. We asked participants with a healthcare background to imagine they were developing a model to predict patients with high future “medical acuity,” and we asked participants with an education background to imagine they were developing a model to predict which students were “academically at-risk” (see Appendix A). [^2] P1 was shown the healthcare vignette due to data availability constraints encountered at early phases of the interview schedule. During the interview, P1 signaled places where their interpretation of the vignette was informed by their domain experience. We took this into consideration during analysis. We provided participants a set of features and several outcome variables and instructed them to imagine that these were the data available for model development.
During the vignette, we asked participants to reflect on which outcomes might serve as a “better or worse choice” of prediction target given the stated modeling goals. After learning how participants initially approached this task, we then presented a series of model evaluation plots indicating the performance of predictive models targeting each outcome (see Appendix A). After showing each plot, we paused and asked participants to again reflect on which outcomes might serve as better or worse choices for the stated modeling goals. We also asked participants to share any additional information that they would ideally like to have, to better inform their thought process. As participants engaged with the task over subsequent interviews, we refined the specificity of the scenario description and accompanying evaluation plots to further probe issues that surfaced in prior interviews. This second segment lasted 20-30 minutes on average.
@@ 3.2. Participants
Participants were recruited via purposive sampling techniques (Heckathorn, 2011) from authors’ professional networks and social media (Table 1). To ensure that participants had a baseline level of domain expertise, we required that data scientists had domain-specific graduate-level training related to healthcare or education (e.g., biostatistics, learning sciences) or at least one year of full-time industry experience in a relevant domain to be eligible to participate. We further required that participants had prior experience developing and evaluating predictive models as part of their role. Our participants had a range of post-graduate experience at the time of the interview, spanning from 3-6 months to 15 years. At the time of the interview, participants had a variety of primary roles (e.g., Data Scientist, Assistant Professor, Board Member, PostDoc) from both academic and industry affiliations. As a result, our sample consists of data scientists with a baseline level of domain knowledge and a range of technical experience. This study was approved by the University Institutional Review Board and all participants were compensated with a \$35 Amazon gift card for completing the study.
@@ 3.3. Data and Analysis
One of the authors conducted interviews remotely via Zoom between September and December 2023. In total, we recorded and transcribed 14 hours and 30 minutes of audio after collecting verbal and written consent from all participants. Two authors independently performed open coding on transcripts and met regularly to discuss and reconcile interpretation differences. After interviewing fifteen participants, we observed a common set of themes recurring across interviews and took this as an indicator that we had reached saturation (Saunders et al., 2018). We adopted a reflexive thematic analysis approach (Braun and Clarke, 2006; Clarke and Braun, 2017) to group codes into successively higher themes, informed by discussions among all of the authors. Our process gave rise to sixteen high-level themes, organized into three categories: (1) factors driving the target variable construction process, which we discuss in $\S$§4.1; (2) data scientists’ model development and evaluation practices, which we discuss in $\S$§4.2 and $\S$§4.3; and (3) data scientists’ orientation towards the problem formulation process, which we discuss throughout the findings, when drawing connections to the concept of bricolage.
@@ 3.4. Positionality
Figure 2. An illustration of the target variable construction process presented in our findings. Data scientists specify an initial prediction task based on their available data, then iteratively refine their prediction task by applying (re)formulation strategies. Data scientists proceed with their final prediction task if it satisfies all criteria, or discontinue their project if strategies are exhausted.
@@ 4. Findings
In the following subsections, we detail this bricolage process by describing how data scientists balance validity with other competing criteria ($\S$§ 4.1), the (re)formulation strategies they apply to navigate this balancing act ($\S$§ 4.2), and the theory-driven and data-driven approaches they use to evaluate the validity of their formulation ($\S$§ 4.3).
Figure 3. Data scientists evaluated the validity, predictive performance, portability, simplicity, and resource requirements of their problem formulation.
@@
Predictive Performance The extent to which a model’s predictions align with observed outcomes (e.g., accuracy or AU-ROC). Predicting cardiovascular rather than renal diabetes health outcomes (P10). See $4.2.3\S\ref${subsec:swapping}§.
Resource Requirements The organizational resources required to implement and maintain the formulation. Making do with limited test scores to avoid imposing additional testing (P6). See $\S$§ 4.1.2.
@@
Simplicity The complexity of a task definition or model’s structure and interpretation. Switching from a complex multi-outcome model due to challenges in explanation (P2). See $\S$§ 4.2.2.
Table 2. A description of criteria that data scientists considered while evaluating their problem formulation and illustrative examples of how participants balanced each with validity.
@@ 4.1. Balancing Validity With Other Criteria
While engaging in target variable construction, data scientists were forced to make do with the data they already had available for the task at hand. This process of making do forced data scientists to navigate a rich space of tradeoffs between validity and other important criteria, such as predictive performance, simplicity, resource requirements, and the ease by which it could be translated to different geographic or institutional contexts (i.e., portability) (Figure 3). In Table 2, we provide an overview of illustrative tradeoffs that data scientists made between validity and other criteria. We provide a more detailed description of how the process of making tradeoffs unfolded in the sections below.
@@ 4.1.1. Balancing Validity with Simplicity
While constructing prediction tasks, several data scientists navigated a tension between the validity and simplicity of their formulation. In particular, data scientists found complex formulations, such as those combining multiple outcome variables, hard to explain to other stakeholders and maintain in production. For example, P1 was a director of a data science team at a large education company. Because the essay scoring tool designed by their team was used by schools throughout the country, P1 described the simplicity of their formulation as critical for ensuring the system was robust and maintainable. P1 had considered adopting a sophisticated multi-target formulation to improve validity. Yet, their team later abandoned the approach over practical concerns, explaining that the sophisticated models “are a real hassle in practice. The extra […] bells and whistles […] made it flakier code, something that we didn’t necessarily want to deploy at scale, compared to a relatively straightforward single output variable model.” Further, P1 later shared that they felt it was possible to proceed with using a simple single output variable model if they made several hyperparameter adjustments and user interface tweaks. In this experience, P1 was willing to settle for a specification that had minor validity drawbacks because they felt it would be simple and easy to maintain at scale.
@@ 4.1.2. Balancing Validity with Resource Requirements
Data scientists also evaluated the validity of their formulation against its resource requirements. For example, P6 was a data scientist developing an online math tutoring platform. P6 explained that the time and cost involved with acquiring “gold standard” student test scores to evaluate their models was a frequent challenge. School districts often had data sharing restrictions. Opportunities to test students were limited. However, P6 explained that even if they could test students more often, they would not recommend that approach. Testing students is “just removing opportunity. Removing time from the students’ school calendar, which is already […] crowded.” Instead of collecting data, P6 devised a creative workaround that achieved their project’s goals using the data they already had available. [^3] We detail the strategy P6 and their team used in Section 4.2.4. While P6 acknowledged that their workaround was imperfect, they were fine with proceeding so long as they had “at least some confidence” in its validity. In particular, while reflecting on potential drawbacks of their modeling decisions, P6 shared that “this [application] isn’t high stakes” when compared to other potential modeling contexts (e.g., university admissions). However, had the model been deployed in other contexts, they explained that they might apply a different level of stringency to the validity of their approach.
@@ 4.1.3. Elasticity of Evaluation Criteria
We observed that data scientists applied varying levels of flexibility to different evaluation criteria, a characteristic we term elasticity. Participants frequently demonstrated elasticity with validity standards, tolerating imperfections when doing so reduced the complexity or resource requirements of their formulation. Data scientists actively calibrated this tolerance based on the success of compensatory adjustments (e.g., user interface tweaks) and the perceived “stakes” of their modeling decisions. Interestingly, participants maintained more rigid standards for other criteria. For instance, several participants cited hard cutoffs for acceptable predictive performance — e.g., P8 applied an AU-ROC threshold of 0.7 to determine whether to proceed with a project. This rigidity in predictive performance evaluation likely stems from the ease with which it can be distilled into a single, quantifiable metric, unlike validity, which requires a more holistic assessment involving multiple sources of evidence (Jacobs and Wallach, 2021). While tradeoffs are an inherent component of bricolage, motivated by adaptive response to resource constraints, the process of navigating tradeoffs remains poorly supported. We describe opportunities to help data scientists navigate tradeoffs in Section 5.3.1.
@@ 4.2. Strategies for (Re)formulating Prediction Tasks
Given their limited and imperfect data resources, how do data scientists engage in bricolage to craft a formulation that satisfies their objectives? Much like a painter refining a painting through successive brushstrokes, or a programmer compiling and re-compiling a system after adjusting lines of code (Turkle, 2011), data scientists iteratively construct target variables through (re)formulation strategies. These strategies form the backbone of the dialogue between the data scientist and their formulation, enabling them to respond adaptively to constraints encountered while working with existing data. At times, data scientists apply these strategies instinctively, without a second thought. At other times, they shared eureka moments they encountered after discovering a strategy that leveraged unexpected data sources or combined information in novel ways. In the following sections, we detail five key strategies—piggybacking, composing, swapping, bridging, and refining—that data scientists used to navigate the constraints of their limited repertoire.
@@ 4.2.1. Piggybacking
During the initial stages of the problem formulation process, data scientists often constructed target variables through piggybacking: adopting or extending a problem formulation that was previously used in a similar context in the past. Participants justified piggybacking by citing precedents from related projects. For example, while developing a model to inform university admissions decisions, P5 suggested predicting whether students’ college GPA exceeded a 3.0 because other data science teams used it for similar admissions pipelines in the past. Piggybacking provided a rationale for P5’s formulation decisions and offered a point of comparison with prior work. However, they were also wary that piggybacking could introduce unintended validity consequences.
@@
Piggybacking off of other modeling contexts did not work for P1’s dataset - optimizing for Quadratic Weighted Kappa yielded a model that almost never gave a perfect score. This led P1 to eventually adopt a performance metric better suited to their setting, which balanced predictive accuracy with some chance that students could get a perfect score. P1’s experience illustrates that top-down, theoretical analysis of an established precedent (Quadratic Weighted Kappa) was insufficient for making a well-informed piggybacking decision. Instead, P1’s experience illustrates the “science of the concrete” characteristic of bricolage, where tangible interactions with materials are important for understanding how to assemble elements in the repertoire into a functional artifact.

### v2

@@ 2.5. Measurement and Validity in AI
A growing line of research traces AI system failures to validity pitfalls (Jacobs and Wallach, 2021; Milli et al., 2021; Stray et al., 2021; Guerdan et al., 2023b; Goyal et al., 2022; Obermeyer et al., 2019; Mullainathan and Obermeyer, 2021; Coston et al., 2023). In the quantitative social sciences, the validity of a measurement instrument is established via a multifaceted assessment along several dimensions. Construct validity describes the extent to which a measurement instrument accurately reflects the theoretical construct it purports to measure (Jacobs and Wallach, 2021). In predictive modeling, construct validity is most often discussed in connection to whether the target variable predicted by a model reflects a theoretical construct of interest to model developers (Guerdan et al., 2023b; Coston et al., 2023).  [^1] Jacobs and Wallach (2021) provide a detailed review of sub-dimensions of construct validity as they pertain to measuring fairness properties of algorithmic systems. Internal and external validity are also important dimensions in experimental research. Internal validity describes the extent to which an experimental study can confidently establish a causal relationship between independent and dependent variables (Gergle and Tan, 2014). In predictive modeling, it relates to the existence of a defensible causal relationship between features and a target variable (Coston et al., 2023). External validity, on the other hand, measures how well an analysis can be generalized beyond the specific conditions of the study. This dimension has been linked to the extent to which a predictive model’s in-distribution performance generalizes across real-world deployment contexts (Coston et al., 2023).
@@
We expand upon Mussgnug (2022)’s argument by characterizing how data scientists weigh validity and predictive performance during problem formulation. In support of Mussgnug (2022)’s argument, we find that many prediction tasks pursued by data scientists are indeed measurement problems. For example, some participants in our study used supervised learning models to measure concepts recorded at future points in time — e.g., “10-year cardiovascular risk.” Others used supervised learning models as measurement instruments while measuring a concept at a current point of time — e.g., to score the “authenticity” of a student’s writing. In both cases, data scientists heavily anchored on predictive performance while evaluating their problem formulation. However, counter to Mussgnug (2022)’s argument, we find that data scientists do also engage with validity concepts during problem formulation, albeit with a more pragmatic, situated approach that lacks formal terminology. We uncover how data scientists evaluate validity in our findings ($\S$ 4), and unpack implications for data science pedagogy ($\S$ 5.2.3) and tooling support ($\S$ 5.3.4) in the discussion.
@@ 3. Methods
In this study, we interviewed fifteen data scientists from the education (N=8) and healthcare (N=7) sectors who had prior experience with developing and evaluating predictive models. We conducted one-hour semi-structured interviews to understand participants’ current practices, perceived challenges, and future opportunities for target variable construction in predictive modeling tasks.
@@ 3.1.1. Directed Storytelling
During the first interview segment, we used a directed storytelling approach (Evenson, 2016), inviting data scientists to reflect on a time in the past when they developed and evaluated a predictive model. To ground our discussions, we first asked participants to share a few concrete experiences where they had developed and/or evaluated predictive models as part of their role. Building on these experiences, the interviewer then narrowed in on one of the projects mentioned and asked participants the following follow-up questions, in a semi-structured fashion:
@@
We also asked participants additional follow-up questions when these were not already covered by participants during their earlier responses, such as measurement guidelines and best practices they consulted throughout their project. On average, the first segment lasted 30-40 minutes.
@@ 3.1.2. Vignette-Based Problem Formulation Task
During the second segment, we asked participants to reason about a hypothetical problem formulation as we walked them through a vignette. We aimed to better understand how participants reason about measurement decisions on-the-fly, while potentially surfacing challenges they might not have recalled while reflecting on their prior experiences. We asked participants with a healthcare background to imagine they were developing a model to predict patients with high future “medical acuity,” and we asked participants with an education background to imagine they were developing a model to predict which students were “academically at-risk” (see Appendix A). [^2] P1 was shown the healthcare vignette due to data availability constraints encountered at early phases of the interview schedule. During the interview, P1 signaled places where their interpretation of the vignette was informed by their domain experience. We took this into consideration during analysis. We provided participants a set of features and outcome variables and instructed them to imagine that these were the data available for model development.
During the vignette, we asked participants to reflect on which outcomes might serve as a “better or worse choice” of prediction target given the stated modeling goals. After learning how participants initially approached this task, we then presented a series of model evaluation plots indicating the performance of predictive models targeting each outcome (see Appendix A). After showing each plot, we paused and asked participants to again reflect on which outcomes might serve as better or worse choices for the stated modeling goals. We also asked participants to share any additional information that they would ideally like to have, to better inform their thought process. As participants engaged with the task over subsequent interviews, we refined the specificity of the scenario description and accompanying evaluation plots to further probe issues that surfaced in prior interviews. This second segment lasted 20-30 minutes of the interview on average.
@@ 3.2. Participants
Participants were recruited via purposive sampling techniques (Heckathorn, 2011) from authors’ professional networks (Table 1). To ensure that participants had a baseline level of domain expertise, we required that data scientists had domain-specific graduate-level training related to healthcare or education (e.g., biostatistics, learning sciences) or at least one year of full-time industry experience in a relevant domain to be eligible to participate. We further required that participants had prior experience developing and evaluating predictive models as part of their role. Our participants had a range of post-graduate experience at the time of the interview, spanning from 3-6 months to 15 years. At the time of the interview, participants had a variety of primary roles (e.g., Data Scientist, Assistant Professor, Board Member, PostDoc) from both academic and industry affiliations. As a result, our sample consists of data scientists with a baseline level of domain knowledge and a range of technical experience levels. This study was approved by the University Institutional Review Board and all participants were compensated with a \$35 Amazon gift card for completing the study.
@@ 3.3. Data and Analysis
One of the authors conducted interviews remotely via Zoom between September and December 2023. In total, we recorded and transcribed fourteen hours and thirty minutes of audio after collecting verbal and written consent from all participants. Two authors independently performed open coding on transcripts and met regularly to discuss and reconcile interpretation differences. After interviewing fifteen participants, we observed a common set of themes recurring across interviews and took this as an indicator that we had reached saturation (Saunders et al., 2018). We adopted a reflexive thematic analysis approach (Braun and Clarke, 2006; Clarke and Braun, 2017) to group codes into successively higher themes, informed by synchronous discussions among all of the authors. Our process gave rise to sixteen high-level themes, organized into three categories: (1) factors driving the target variable construction process, which we discuss in $\S$4.1; (2) data scientists’ model development and evaluation practices, which we discuss in $\S$4.2 and $\S$4.3; and (3) data scientists’ orientation towards the problem formulation process, which we discuss throughout the findings, when drawing connections to the concept of bricolage.
@@ 3.4. Positionality
Figure 2. An illustration of the target variable construction process presented in our findings. During target variable construction, data scientists specify an initial prediction task based on their available data, then iteratively refine their prediction task by applying (re)formulation strategies. Data scientists proceed with their final prediction task if it satisfies all criteria, or discontinue their project if strategies are exhausted.
@@ 4. Findings
In the following subsections, we detail this bricolage process by describing how data scientists balance validity with other competing criteria ($\S$ 4.1), apply (re)formulation strategies ($\S$ 4.2), and the engage with theory-driven and data-driven evaluation activities to evaluate validity ($\S$ 4.3).
Figure 3. The five criteria data scientists evaluated during target variable construction.
@@
[…]
Resource Requirements The organizational resources required to implement and maintain the formulation. Making do with limited test scores to avoid imposing additional testing (P6). See $\S$ 4.1.2.
@@
Simplicity The complexity of a task definition or model’s structure and interpretation. Switching from a complex multi-outcome model due to challenges in explanation (P2). See $\S$ 4.2.2.
Table 2. A description of criteria that data scientists considered while evaluating their problem formulation. Right column provides an illustrative example of how participants balanced each criterion with validity.
@@ 4.1. Balancing Validity With Other Criteria
While engaging in target variable construction, data scientists were forced to make do with the data they already had available for the task at hand. This process of making do forced data scientists to navigate a rich space of tradeoffs between validity and other important criteria, such as predictive performance, simplicity, resource requirements, and the ease by which it could be translated to different geographic or institutional contexts (i.e., portability) (Figure 3). In Table 2, we provide an overview of illustrative tradeoffs that data scientists made between validity and other criteria. We provide a description of how this process of making tradeoffs unfolded in the sections below.
@@ 4.1.1. Balancing Validity with Simplicity
While constructing prediction tasks, several data scientists navigated a tension between the validity and simplicity of their formulation. In particular, data scientists found complex formulations, such as those combining multiple outcome variables, hard to explain to other stakeholders and maintain in production. For example, P1 was a director of a data science team at a large education company. Because the essay scoring tool designed by their team was used by schools throughout the country, P1 described the simplicity of their formulation as critical for ensuring the system was robust and maintainable. P1 had considered adopting a sophisticated multi-target formulation to improve validity. Yet, their team later abandoned the approach over practical concerns, explaining that the sophisticated models “are a real hassle in practice. The extra […] bells and whistles […] made it flakier code, something that we didn’t necessarily want to deploy at scale, compared to a relatively straightforward single output variable model.” Further, P1 later shared that they felt it was possible to proceed with using a simple single output variable model if they made several hyperparameter adjustments and user interface tweaks. In this experience, P1 was willing to settle for a problem formulation that had minor validity drawbacks because they felt it would be simple and easy to maintain in a production system.
@@ 4.1.2. Balancing Validity with Resource Requirements
Data scientists also evaluated the validity of their formulation against its resource requirements. For example, P6 was a data scientist developing an online math tutoring platform. P6 explained that the time and cost involved with acquiring “gold standard” student test scores to evaluate their models was a frequent challenge. School districts often had data sharing restrictions. Opportunities to test students were limited. However, P6 explained that even if they could test students more often, they would not recommend that approach. Testing students is “just removing opportunity. Removing time from the students’ school calendar, which is already […] crowded.” Instead of collecting data, P6 devised a creative workaround that achieved their project’s goals using the data they already had available. [^3] We detail the strategy P6 and their team used in Section 4.2.4. While P6 acknowledged that their workaround was imperfect, they were fine with proceeding so long as they had “at least some confidence” in its validity. In particular, while reflecting on potential drawbacks of their modeling decisions, P6 shared that “this [application] isn’t high stakes” when compared to other potential modeling contexts (e.g., university admissions). They explained that had the model been deployed in other contexts, they might apply a different level of stringency to the validity of their approach.
@@ 4.1.3. Elasticity of Evaluation Criteria
Data scientists applied varying levels of flexibility to different evaluation criteria, a characteristic we term elasticity. Participants frequently demonstrated elasticity with validity standards, tolerating imperfections when doing so reduced the complexity or resource requirements of their formulation. Data scientists calibrated this tolerance based on the success of compensatory adjustments (e.g., user interface tweaks) and the perceived “stakes” of their modeling decisions. Interestingly, participants maintained more rigid standards for other criteria. For instance, several participants cited hard cutoffs for acceptable predictive performance — e.g., P8 applied an AU-ROC threshold of 0.7 to determine whether to proceed with a project. This rigidity in predictive performance evaluation likely stems from the ease with which it can be distilled into a single, quantifiable metric, unlike validity, which requires a more holistic assessment involving multiple sources of evidence (Jacobs and Wallach, 2021). While tradeoffs are an inherent component of bricolage, motivated by adaptive response to resource constraints, the process of navigating tradeoffs remains poorly supported. We describe opportunities to help data scientists navigate tradeoffs in Section 5.3.1.
@@ 4.2. Strategies for (Re)formulating Prediction Tasks
Given their limited and imperfect data resources, how do data scientists engage in bricolage to craft a formulation that satisfies their criteria? Much like a painter refining a painting through successive brushstrokes, or a programmer compiling and re-compiling a system after adjusting lines of code (Turkle, 2011), data scientists iteratively construct target variables through (re)formulation strategies. These strategies form the backbone of the dialogue between the data scientist and their formulation, enabling them to respond adaptively to constraints encountered while working with existing data. At times, data scientists apply these strategies instinctively, without a second thought. At other times, encountered eureka moments after discovering a strategy that leveraged unexpected data sources or combined existing ones in novel ways. In the following sections, we detail five key strategies—piggybacking, composing, swapping, bridging, and refining—that data scientists used to navigate the constraints of their limited repertoire.
@@ 4.2.1. Piggybacking
During the initial stages of problem formulation, data scientists often constructed target variables through piggybacking: adopting or extending a problem formulation that was previously used in a similar context in the past. Participants justified piggybacking by citing precedents from related projects. For example, while developing a model to inform university admissions decisions, P5 suggested predicting whether students’ college GPA exceeded a 3.0 because other data science teams used it for similar admissions pipelines in the past. Piggybacking provided a rationale for P5’s formulation decisions and offered a point of comparison with prior work. However, they were also wary that piggybacking could introduce unintended validity consequences.
@@
Piggybacking off of other modeling contexts did not work for P1 - optimizing for Quadratic Weighted Kappa yielded a model that almost never gave a perfect score. This led P1 to eventually adopt a performance metric better suited to their setting, which balanced predictive accuracy with some chance that students could get a perfect score. P1’s experience illustrates that top-down, theoretical analysis of an established precedent (Quadratic Weighted Kappa) was insufficient for making a well-informed piggybacking decision. Instead, P1’s experience illustrates the “science of the concrete” characteristic of bricolage, where tangible interactions with materials are important for understanding how to assemble elements in the repertoire into a functional artifact.